Central Park, which lies in the middle of Manhattan, is a rectangular-shaped and has a 6-mile perimeter. The length is 5 times the width. What are the dimensions of the park?
Answer by rfer(16322)

2x+2(5x)=6
2x+10x=6
12x=6
x=6/12=.5 mi
5x=2.5 mi